# Beyond Preferences in AI Alignment

### v1

@@ 1 Introduction
Of course, preferentism in AI alignment is not without its critics. Over the years, there has been considerable discussion as to whether its component theses are warranted (Shah,, 2018; Eckersley,, 2018; Hadfield-Menell and Hadfield,, 2018; Wentworth,, 2019; Gabriel,, 2020; Vamplew et al.,, 2022; Korinek and Balwit,, 2022; Thornley,, 2023), echoing similar debates in economics, decision theory, and philosophy. Nonetheless, it is apparent that the dominant practice of AI alignment has yet to absorb the thrust of these debates. Consequently, we believe it is worthwhile to identify the descriptive and normative commitments of preferentist approaches, to state clearly their limitations, and to describe conceptual and technical alternatives that are ripe for further research (Table 1).
@@ 2 Beyond rational choice theory when modeling humans
Rational choice theory in machine learning. In keeping with rational choice theory, many machine learning and AI systems also assume that human preferences can be derived from human choices in a more or less direct manner, and furthermore represent those preferences in terms of scalar utilities or rewards. This is most pronounced in the fields of inverse reinforcement learning (Ng and Russell,, 2000; Abbeel and Ng,, 2004; Hadfield-Menell et al.,, 2016) and reinforcement learning from human feedback (Christiano et al.,, 2017; Zhu et al.,, 2023), which explicitly assume that the behavior of a human can be described as (approximately) maximizing a sum of scalar rewards over time, and then tries to infer a reward function that explains the observed behavior. Similar assumptions can be found in the field of recommender systems (Thorburn et al.,, 2022), with many papers modeling recommendation as the problem of showing items to users that they are most likely to engage with, which is presumed to be the item they find the most rewarding (Li et al.,, 2010; Hill et al.,, 2017; McInerney et al.,, 2018).
@@ 2.1 Beyond noisily-rational models of human decisions
Challenges to modeling bounded rationality. How might AI systems that infer human preferences and values account for these findings? One approach might be to incorporate a sufficiently long list of known heuristics and biases into our models of human decision-making, thereby ensuring that preferences can be robustly inferred even in the presence of such biases (Evans et al.,, 2016; Chan et al.,, 2021). However, this approach is highly contingent upon on our current state of knowledge about human rationality — what if we miss out important biases in our models, leading to inaccurate predictions and inferences? (Christiano, 2015b, ; Steinhardt,, 2017) As a potential remedy, Shah et al., (2019) suggest learning human biases alongside their preferences. But a conceptual difficulty remains: Without any inductive constraints on the types of errors humans are susceptible to, how can we ensure that human biases are accurately learned? As Armstrong and Mindermann, (2018) show, even inductive preferences for more parsimonious models of human decision-making cannot distinguish important classes of observationally-equivalent hypotheses from intuitively plausible ones, such as the possibility that humans are acting anti-rationally by minimizing the satisfaction of their preferences.
@@ 2.2 Beyond reward and utility functions as representations of human preferences
The limited expressivity of reward functions. These issues are most easily appreciated in the case of (scalar, Markovian) reward functions. As noted earlier, the reward representation assumes that the utility of a sequence of states and actions $\xi=(s_{1},a_{1},...,a_{T-1},s_{T})$ can be decomposed into a sum of scalar rewards over time:
@@ 2.3 Beyond preferences as representations of human values and reasons
Preferences are constructed, not basic. Thus far, we have proceeded as if human motivations and values are adequately captured by the concept of “preference” as it is used in rational choice theory. But as far as evaluative concepts go, this concept of “preference” is an extremely thin one: Mathematically, a “preference” is just some ordering of two options, which can be interpreted as either a disposition to choose one option over another, subjective liking of one option over the other Franklin et al., (2022), or an all-things-considered judgment in favor of one of the options. Distinct as these interpretations are, what they share is their highly abstract and general nature — “preference” is a thin concept because it does not encode richer semantic information beyond the bare notion of “betterness”. Insofar as utility functions are interpreted as representations of preferences, this thinness is inherited by them: Utility just represents the mere preferability of some option. But why exactly are some options preferred over others? In virtue of what reasons do people make these preference judgments? Without answering these questions, we are unlikely to model how someone’s preferences generalize to novel options in ways they would endorse. To do so, we must go beyond preferences as the fundamental unit of analysis, and understand how preferences are computed and constructed from our reasons and values (Warren et al.,, 2011; Lichtenstein and Slovic,, 2006).
@@
Utility functions as aggregators of distinct evaluative judgments. How should AI systems model such evaluative concepts, and their relationship to preferences and action? As a first pass, one might turn the utility representation theorems on their head, viewing reward and utility functions as generators of human preferences, instead of mere representations of them. Indeed, as gestured at earlier, reward and utility functions are often interpreted in this way, with rewards, costs, and utilities respectively treated as biological signals (Singh et al.,, 2009), energetic expenditure (Ab Azar et al.,, 2020), or units of pleasure (Bentham,, 1789). Preferences can then be treated as downstream comparisons of these more basic quantities, as assumed in reinforcement learning from human feedback (Christiano et al.,, 2017; Knox et al.,, 2024; Zhu et al.,, 2023). Taking this line of thought further, one might treat evaluative concepts such as “aesthetic quality” or “helpfulness” as features over which a reward or utility function is defined, reducing the problem of “value learning” to one of representation or feature learning (Barreto et al.,, 2017; Bobu et al.,, 2022, 2024). On this interpretation, reward and utility functions represent aggregate evaluative judgments, with each feature corresponding to a distinct way of valuing the world.
Utility functions assume that values are always commensurable. Although there is much to be said in favor of this approach, we believe that it is not quite enough. For one, it is still subject to the representational limits of reward and utility functions. In particular, if utility functions are used to represent aggregate value judgments, this effectively assumes that distinct human values are always commensurable in some way, and that our resulting preferences are always complete. Yet, as value pluralists argue, there are contexts where it seems hard or impossible to commensurate our values (Anderson,, 1995), resulting in choices where our reasons run short, and we cannot say if one option is ultimately better than another (Chang,, 1997). [^10] See our immigration example from earlier, where it may be unclear how to prioritize between political freedom and material security when deciding whether to migrate. Even when we do commensurate our values, utility function do not provide further information on our reasons and justifications for those trade-offs.
@@ 3.1 Beyond expected utility theory as an analytical lens
Coherence is not rationally required. However, coherence arguments for expected utility theory are not as strong as the AI alignment literature has often presumed. The most extensive version of these arguments is given by Gustafsson, (2022), who provides a money pump argument for preference completeness, and then uses completeness to derive arguments for transitivity, continuity, and independence. Yet, as Thornley, (2023) points out, the argument for completeness depends on particular assumptions about how agents are permitted to choose when offered a series of potentially exploitative trades, which can be avoided as long as agents do not accept offers that are less preferred than options they previously turned down. [^15] Note that whereas Gustafsson, (2022) is focused on justifying the VNM axioms as requirements of rationality (in part by introducing and arguing for other principles of rationality, such as Decision-Tree Separability), Thornley, (2023) is focused on whether the VNM axioms will apply to advanced AI systems, and takes no position on whether they are rationally required. Here we go one step further, and suggest that arguments by Thornley, (2023) and Petersen, (2023) place strong pressure on Gustafsson’s acceptance of rationality principles like Decision-Tree Separability, and hence the argument that the VNM axioms are rationally required. Petersen, (2023) formalizes this counter-argument further, proposing a dynamic choice rule that ensures agents with incomplete preferences are invulnerable to money pumps. [^16] Analogous arguments have made in defense of imprecise probabilities (Bradley and Steele,, 2014), since they imply incomplete preferences over uncertain outcomes. See also Laibson and Yariv, (2007) for an argument that money-pumping traders are eliminated in competitive markets, thereby protecting non-coherent preferences. Indeed, it is accepted by many decision theorists that preference completeness is not a requirement of rationality; instead, all that is required is for an agent’s preferences to be coherently extendible (Steele and Stefánsson,, 2020). In turn, this implies that rational agents need not be representable as EU maximizers.
@@
Alternative analytical lenses to EUT. What alternatives might one turn to instead to ground understanding, prediction, and alignment of advanced AI systems? Since many others have already addressed some version of these questions question, we offer here a brief taxonomy of approaches.
@@ 3.2 Beyond globally coherent agents as design targets
Maintaining locality despite global scope. Suppose, however, that some actors want to build advanced AI systems that are not bounded in these ways. For example, many AI companies are keen to develop general purpose AI assistants, which follow human instructions in a wide range of domains and contexts, remain operational across contexts, and possess enough understanding of the wider world that they can represent both themselves and their users as entities in that world model. LLMs are increasingly being used in this way, and while their reasoning capabilities remain unreliable and limited (Valmeekam et al., 2023b, ; Dziri et al.,, 2023; Momennejad et al.,, 2024), one might imagine augmenting or embedding them within systems with more coherent representations and reasoning abilities (Parisi et al.,, 2022; Sumers et al.,, 2023). Can we ensure that such systems continue to function as tools, despite their increasingly global scope?
@@ 3.3 Beyond preferences as the normative basis of action
The need for theories of normative reasoning. What options do we have for doing this? What would it look like to reason about the preferences and values one ought to have? Given the complexity of these questions, one might hope to sidestep the need for a formal account like EUT entirely, and instead train AI systems to imitate human normative reasoning. This is exemplified by the standard training objective of LLMs, which incentivizes replication of human-generated text. By training such systems on normative human judgments, one might hope that LLMs will learn the reasoning patterns that produce such judgments (Jiang et al.,, 2021). Recent methods such as Constitutional AI (Bai et al., 2022b, ) take this idea one step further, bootstrapping an LLM’s ability to approximate human normative judgments by generating self-critiques (Saunders et al.,, 2022) and revisions, then finetuning the LLM on its own revisions. However, even strong LLMs currently struggle to reproduce human judgments on sufficiently nuanced normative questions (Jin et al.,, 2022; Kwon et al., 2023a, ), and there are reasons to doubt whether LLMs can learn to reliably reason through either imitation (van Rooij et al.,, 2023; Dziri et al.,, 2023) or self-critique (Stechly et al.,, 2023; Valmeekam et al., 2023a, ). This unreliability suggests that we might want formal theories of normative reasoning after all. Without such theories, we would have no general way of evaluating whether an AI system reasons “correctly”, beyond comparison to often fallible human judgments. [^26] While formal theories of reasoning will ultimately have to be evaluated against human judgments themselves, they deliver systematicity and precision that many AI systems do not. Just as with mathematics, logic, and probability theory, formal reasoning systems can succinctly express what we would reflectively endorse, provided that we accept certain principles of reasoning as sound. Perhaps imitation or self-critique will be enough for the majority of everyday situations, but if we want AI systems to address normative questions that are increasingly far afield from past human experience, the ability to validate or produce long chains of normative reasoning may be crucial for both system evaluation and scalable oversight.
@@ 4 Beyond single-principal AI alignment as preference matching
Foundational limitations of reward learning. For all its success, RLHF faces numerous technical challenges (Casper et al.,, 2023), ranging from issues with preference elicitation (Knox et al.,, 2023) and scalable oversight (Leike et al.,, 2018) to over-optimization (Gao et al.,, 2023; Moskovitz et al.,, 2024) and stable training (Hejna et al.,, 2024). Our focus, however, is more foundational, and applies to not just RLHF but any alignment method derived from reward learning: [^28] This includes Direct Preference Optimization (Rafailov et al.,, 2024), Contrastive Preference Learning (Hejna et al.,, 2024), and Distributional Preference Learning (Siththaranjan et al.,, 2024). By committing to a reward representation of human preferences or values, reward learning suffers from all the representational limits we discussed in Section 2. Furthermore, by treating reward as something to be optimized, reward-based methods adopt EUT as a normative standard, with all the issues that Section 3 describes.
@@ 4.1 Beyond alignment with scalar and acontexual rewards
Context-sensitive preference models as an intermediate solution. What would it take to align an AI system that operates across contexts? One option is the use of context-sensitive reward functions, as described in Section 3. Though this approach runs the risk of context-manipulating incentives, it may well be adequate for sufficiently bounded systems. Similar to our ECD proposal in Section 2, context-sensitivity could be achieved by per context commensuration of multiple values, perhaps by learning separate reward or preference models for each value (Wu et al.,, 2024; Go et al.,, 2024), then aggregating their rewards with different weights depending on the downstream context. Context switches could then be triggered by users by switching between their desired “mode”, as has been implemented in some LLM-based AI assistants (Edwards,, 2023).
@@ 4.3 Beyond preferences as the target of alignment
Alignment with role-specific normative criteria. This basic point, of course, is not new: As many have long appreciated, identifying someone’s welfare or best interests with their preferences runs into a thicket of philosophical issues (Sen et al.,, 1999; Nussbaum,, 2001). Recognizing these issues, Gabriel, (2020) argues for an explicitly moral conception of alignment: “the agent does what it morally ought to do, as defined by the individual or society”. [^31] Gabriel, (2020) uses “values” to describe this alignment target, though in a slightly narrower sense than ours. Whereas we have primarily used “values” to refer to evaluative concepts and judgments in general, Gabriel’s use implicitly picks out the values that are normatively relevant to AI system behavior. Others have proposed similar approaches, though they replace “morally ought” with what an agent or humanity as a whole would reflectively endorse, as in ideal observer theories (Firth,, 1952; Brandt,, 1955) or coherent extrapolated volition (Yudkowsky,, 2004). However, it is far from clear how to operationalize these abstract principles. To make progress, we suggest a conception of single-principal alignment that is significantly more constrained: When an AI system only serves an individual in performing a particular task or role, it should be aligned with the normative ideals or criteria that appropriate for that role. For narrow systems, this requires task-specific determination of appropriate normative criteria. For general-purpose AI assistants, this implies alignment with the normative ideal of an assistant, rather than alignment to an individual’s preferences, or to human normativity writ large.
@@
How does this ideal address the issues with preference alignment that we have raised? Here are a number of suggestions: First, a good assistant does not presume certainty about a person’s preferences and values (Hadfield-Menell et al.,, 2016). This means maintaining an awareness of their own ignorance, while avoiding unwarranted extrapolation of preferences from one context to another, including Knightian uncertainty about how preferences extrapolate (Dalrymple et al.,, 2024). Second, a good assistant is aware that some choices are hard, and some options may seem incomparable (Chang,, 1997). When helping someone with such a choice, the assistant does not pretend to know which option is better, or try to optimize that person’s life; instead, the assistant respects their autonomy, and empowers them to make the most informed choice possible (Du et al.,, 2020), while ultimately remaining agnostic as to which choice is “best”. Third, a good assistant understands and respects the values of the person they are assisting. This means recognizing that a person’s preferences often derive from their values, which can take priority over their immediate requests and preferences (London and Heidari,, 2023). The assistant also enables those values to grow and change through normatively acceptable forms of exploration, reflection, volition, or even drift, while avoiding manipulating them or restricting them (Ammann,, 2023). Finally, a good assistant, being situated in wider society, respects the preferences and values of others (Kirk et al.,, 2024). When assisting someone who wishes harm out of anger, the assistant might dissuade them from acting against their better nature. When asked to directly harm others, the assistant might refuse.

### v2

@@ 1 Introduction
Of course, preferentism in AI alignment is not without its critics. There has been considerable discussion as to whether its component theses are warranted (Shah,, 2018; Eckersley,, 2018; Hadfield-Menell and Hadfield,, 2018; Wentworth,, 2019, 2023; Gabriel,, 2020; Vamplew et al.,, 2022; Korinek and Balwit,, 2022; Garrabrant,, 2022; Thornley,, 2023), echoing similar debates in economics, decision theory, and philosophy. Nonetheless, it is apparent that the dominant practice of AI alignment has yet to absorb the thrust of these debates. Consequently, we believe it is worthwhile to identify the descriptive and normative commitments of preferentist approaches, to state clearly their limitations, and to describe conceptual and technical alternatives that are ripe for further research (Table 1).
@@ 2 Beyond rational choice theory when modeling humans
Rational choice theory in machine learning. In keeping with rational choice theory, many machine learning and AI systems also assume that human preferences can be derived from human choices in a more or less direct manner, and furthermore represent those preferences in terms of scalar utilities or rewards. This is most pronounced in the fields of inverse reinforcement learning (Ng and Russell,, 2000; Abbeel and Ng,, 2004; Hadfield-Menell et al.,, 2016) and reinforcement learning from human feedback (Christiano et al.,, 2017; Zhu et al.,, 2023), which explicitly assume that the behavior of a human can be described as (approximately) maximizing a sum of scalar rewards over time, and then try to infer a reward function that explains the observed behavior. Similar assumptions can be found in the field of recommender systems (Thorburn et al.,, 2022), with many papers modeling recommendation as the problem of showing items to users that they are most likely to engage with, which is presumed to be the item they find the most rewarding (Li et al.,, 2010; Hill et al.,, 2017; McInerney et al.,, 2018).
@@ 2.1 Beyond noisily-rational models of human decisions
Challenges to modeling bounded rationality. How might AI systems that infer human preferences and values account for these findings? One approach might be to incorporate a sufficiently long list of known heuristics and biases into our models of human decision-making, thereby ensuring that preferences can be robustly inferred even in the presence of such biases (Evans et al.,, 2016; Chan et al.,, 2021). However, this approach is highly contingent upon on our current state of knowledge about human rationality — what if we miss out important biases in our models, leading to inaccurate predictions and inferences (Christiano, 2015b, ; Steinhardt,, 2017)? As a potential remedy, Shah et al., (2019) suggest learning human biases alongside their preferences. But a conceptual difficulty remains: Without any inductive constraints on the types of errors humans are susceptible to, how can we ensure that human biases are accurately learned? As Armstrong and Mindermann, (2018) show, even inductive preferences for more parsimonious models of human decision-making cannot distinguish intuitively plausible hypotheses from observationally-equivalent but implausible hypotheses, such as the possibility that humans are acting anti-rationally by minimizing the satisfaction of their preferences.
@@ 2.2 Beyond reward and utility functions as representations of human preferences
The limited expressivity of reward functions. These issues are most easily appreciated in the case of (scalar, Markovian) reward functions. As noted earlier, the reward representation assumes that the utility of a sequence of states and actions $\xi=(s_{1},a_{1},...,s_{T},a_{T})$ can be decomposed into a sum of scalar rewards over time:
@@ 2.3 Beyond preferences as representations of human values and reasons
Preferences are constructed, not basic. Thus far, we have proceeded as if human motivations and values are adequately captured by the concept of “preference” as it is used in rational choice theory. But as far as evaluative concepts go, this concept of “preference” is an extremely thin one: Mathematically, a “preference” is just some ordering of two options, which can be interpreted as either a disposition to choose one option over another, subjective liking of one option over the other (Franklin et al.,, 2022), or an all-things-considered judgment in favor of one of the options. Distinct as these interpretations are, what they share is their highly abstract and general nature — “preference” is a thin concept because it does not encode richer semantic information beyond the bare notion of “betterness”. Insofar as utility functions are interpreted as representations of preferences, this thinness is inherited by them: Utility just represents the mere preferability of some option. But why exactly are some options preferred over others? In virtue of what reasons do people make these preference judgments? Without answering these questions, we are unlikely to model how someone’s preferences generalize to novel options in ways they would endorse. To do so, we must go beyond preferences as the fundamental unit of analysis, and understand how preferences are computed and constructed from our reasons and values (Warren et al.,, 2011; Lichtenstein and Slovic,, 2006).
@@
Utility functions as aggregators of distinct evaluative judgments. How should AI systems model such evaluative concepts, and their relationship to preferences and action? As a first pass, one might turn the utility representation theorems on their head, viewing reward and utility functions as generators of human preferences, instead of mere representations of them. Indeed, as gestured at earlier, reward and utility functions are often interpreted in this way, with rewards, costs, and utilities respectively treated as biological signals (Singh et al.,, 2009), energetic expenditure (Ab Azar et al.,, 2020), or units of pleasure (Bentham,, 1789). Preferences can then be treated as downstream comparisons of these more basic quantities, as assumed in reinforcement learning from human feedback (Christiano et al.,, 2017; Knox et al., 2024b, ; Zhu et al.,, 2023). Taking this line of thought further, one might treat evaluative concepts such as “aesthetic quality” or “helpfulness” as features over which a reward or utility function is defined, reducing the problem of “value learning” to one of representation or feature learning (Barreto et al.,, 2017; Bobu et al.,, 2022, 2024). On this interpretation, reward and utility functions represent aggregate evaluative judgments, with each feature corresponding to a distinct way of valuing the world.
Utility functions assume that values are always commensurable. Although there is much to be said in favor of this approach, we believe that it is not quite enough. For one, it is still subject to the representational limits of reward and utility functions. In particular, if utility functions are used to represent aggregate value judgments, this effectively assumes that distinct human values are always commensurable in some way, and that our resulting preferences are always complete. Yet, as value pluralists argue, there are contexts where it seems hard or impossible to commensurate our values (Anderson,, 1995), resulting in choices where our reasons run short, and we cannot say if one option is ultimately better than another (Chang,, 1997). [^10] See our immigration example from earlier, where it may be unclear how to prioritize between political freedom and material security when deciding whether to migrate. Even when we do commensurate our values, utility functions do not provide further information on our reasons and justifications for those trade-offs.
@@ 3.1 Beyond expected utility theory as an analytical lens
Coherence is not rationally required. However, coherence arguments for expected utility theory are not as strong as the AI alignment literature has often presumed. The most extensive version of these arguments is given by Gustafsson, (2022), who provides a money pump argument for preference completeness, and then uses completeness to derive arguments for transitivity, continuity, and independence. Yet, as Thornley, (2023) points out, the argument for completeness depends on particular assumptions about how agents are permitted to choose when offered a series of potentially exploitative trades, which can be avoided as long as agents do not accept offers that are less preferred than options they previously turned down. [^15] Note that whereas Gustafsson, (2022) is focused on justifying the VNM axioms as requirements of rationality (in part by introducing and arguing for other principles of rationality, such as Decision-Tree Separability), Thornley, (2023) is focused on whether the VNM axioms will apply to advanced AI systems, and takes no position on whether they are rationally required. Here we go one step further, and suggest that arguments by Thornley, (2023) and Petersen, (2023) place strong pressure on Gustafsson’s acceptance of rationality principles like Decision-Tree Separability, and hence the argument that the VNM axioms are rationally required. Petersen, (2023) formalizes this counter-argument further, proposing a dynamic choice rule that ensures agents with incomplete preferences are invulnerable to money pumps. [^16] Analogous arguments have made in defense of imprecise probabilities (Bradley and Steele,, 2014), since they imply incomplete preferences. See also Laibson and Yariv, (2007) on how non-EU preferences are protected by competitive markets, and von Widekind, (2008) on how non-EU preferences can be evolutionarily stable. Indeed, it is accepted by many decision theorists that preference completeness is not a requirement of rationality; instead, all that is required is for an agent’s preferences to be coherently extendible (Steele and Stefánsson,, 2020). In turn, this implies that rational agents need not be representable as EU maximizers.
@@
Alternative analytical lenses to EUT. What alternatives might one turn to instead to ground understanding, prediction, and alignment of advanced AI systems? Since many others have already addressed some version of these questions, we offer here a brief taxonomy of approaches.
@@ 3.2 Beyond globally coherent agents as design targets
Maintaining locality despite global scope. Suppose, however, that some actors want to build advanced AI systems that are not bounded in these ways. For example, many AI companies are keen to develop general purpose AI assistants, which follow human instructions in a wide range of domains and contexts, remain operational across contexts, and possess enough understanding of the wider world that they can represent both themselves and their users as entities in that world model. LLMs are increasingly being used in this way, and while their reasoning capabilities remain unreliable and limited (Valmeekam et al., 2023b, ; Dziri et al.,, 2023; Momennejad et al.,, 2024), one might imagine augmenting or embedding them within systems with more coherent representations and reasoning abilities (Parisi et al.,, 2022; Sumers et al.,, 2024). Can we ensure that such systems continue to function as tools, despite their increasingly global scope?
@@ 3.3 Beyond preferences as the normative basis of action
The need for theories of normative reasoning. What options do we have for doing this? What would it look like to reason about the preferences and values one ought to have? Given the complexity of these questions, one might hope to sidestep the need for a formal account like EUT entirely, and instead train AI systems to imitate human normative reasoning. This is exemplified by the standard training objective of LLMs, which incentivizes replication of human-generated text. By training such systems on normative human judgments, one might hope that LLMs will learn the reasoning patterns that produce such judgments (Jiang et al.,, 2021). Recent methods such as Constitutional AI (Bai et al., 2022b, ) take this idea one step further, bootstrapping an LLM’s ability to approximate human normative judgments by generating self-critiques (Saunders et al.,, 2022) and revisions, then finetuning the LLM on its own revisions. However, even strong LLMs currently struggle to reproduce human judgments on sufficiently nuanced normative questions (Jin et al.,, 2022; Kwon et al., 2023a, ), and there are reasons to doubt whether LLMs can learn to reliably reason through either imitation (Van Rooij et al.,, 2024; Dziri et al.,, 2023) or self-critique (Stechly et al.,, 2023; Valmeekam et al., 2023a, ). This unreliability suggests that we might want formal theories of normative reasoning after all. Without such theories, we would have no general way of evaluating whether an AI system reasons “correctly”, beyond comparison to often fallible human judgments. [^26] While formal theories of reasoning will ultimately have to be evaluated against human judgments themselves, they deliver systematicity and precision that many AI systems do not. Just as with mathematics, logic, and probability theory, formal reasoning systems can succinctly express what we would reflectively endorse, provided that we accept certain principles of reasoning as sound. Perhaps imitation or self-critique will be enough for the majority of everyday situations, but if we want AI systems to address normative questions that are increasingly far afield from past human experience, the ability to validate or produce long chains of normative reasoning may be crucial for both system evaluation and scalable oversight.
@@ 4 Beyond single-principal AI alignment as preference matching
Foundational limitations of reward learning. For all its success, RLHF faces numerous technical challenges (Casper et al.,, 2023), ranging from issues with preference elicitation (Knox et al., 2024a, ) and scalable oversight (Leike et al.,, 2018) to over-optimization (Gao et al.,, 2023; Moskovitz et al.,, 2024) and stable training (Hejna et al.,, 2024). Our focus, however, is more foundational, and applies to not just RLHF but any alignment method derived from reward learning: [^28] This includes Direct Preference Optimization (Rafailov et al.,, 2024), Contrastive Preference Learning (Hejna et al.,, 2024), and Distributional Preference Learning (Siththaranjan et al.,, 2024). By committing to a reward representation of human preferences or values, reward learning suffers from all the representational limits we discussed in Section 2. Furthermore, by treating reward as something to be optimized, reward-based methods adopt EUT as a normative standard, with all the issues that Section 3 describes.
@@ 4.1 Beyond alignment with scalar and acontexual rewards
Context-sensitive preference models as an intermediate solution. What would it take to align an AI system that operates across contexts? One option is the use of context-sensitive reward functions (Pitis et al.,, 2024), as described in Section 3. Though this approach runs the risk of context-manipulating incentives, it may well be adequate for sufficiently bounded systems. Similar to our ECD proposal in Section 2, context-sensitivity could be achieved by per context commensuration of multiple values, perhaps by learning separate reward or preference models for each value (Wu et al.,, 2024; Go et al.,, 2024; Xu et al.,, 2024), then aggregating their rewards with different weights depending on the downstream context. Context switches could then be triggered by users by selecting a desired “mode” (Edwards,, 2023) or specifying a system prompt (Pitis et al.,, 2024).
@@ 4.3 Beyond preferences as the target of alignment
Alignment with role-specific normative criteria. This basic point, of course, is not new: As many have long appreciated, identifying someone’s welfare or best interests with their preferences runs into a thicket of philosophical issues (Sen et al.,, 1999; Nussbaum,, 2001). Recognizing these issues, Gabriel, (2020) argues for an explicitly moral conception of alignment: “the agent does what it morally ought to do, as defined by the individual or society”. [^31] Gabriel, (2020) uses “values” to describe this alignment target, though in a slightly narrower sense than ours. Whereas we have primarily used “values” to refer to evaluative concepts and judgments in general, Gabriel’s use implicitly picks out the values that are normatively relevant to AI system behavior. Others have proposed similar approaches, though they replace “morally ought” with what an agent or humanity as a whole would reflectively endorse, as in ideal observer theories (Firth,, 1952; Brandt,, 1955) or coherent extrapolated volition (Yudkowsky,, 2004). However, it is far from clear how to operationalize these abstract principles. To make progress, we suggest a conception of single-principal alignment that is significantly more constrained: When an AI system only serves an individual in performing a particular task or role, it should be aligned with the normative ideals or criteria that are appropriate for that role. For narrow systems, this requires task-specific determination of appropriate normative criteria. For general-purpose AI assistants, this implies alignment with the normative ideal of an assistant, rather than alignment to an individual’s preferences, or to human normativity writ large.
@@
How does this ideal address the issues with preference alignment that we have raised? Here are a number of suggestions: First, a good assistant does not presume certainty about a person’s preferences and values (Hadfield-Menell et al.,, 2016). This means maintaining an awareness of their own ignorance, while avoiding unwarranted extrapolation of preferences from one context to another, including Knightian uncertainty about how preferences extrapolate (Dalrymple et al.,, 2024). Second, a good assistant is aware that some choices are hard, and some options may seem incomparable (Chang,, 1997). When helping someone with such a choice, the assistant does not pretend to know which option is better, or try to optimize that person’s life; instead, the assistant respects their autonomy, and empowers them to make the most informed choice possible (Du et al.,, 2020), while ultimately remaining agnostic as to which choice is “best”. Third, a good assistant understands and respects the values of the person they are assisting. This means recognizing that a person’s preferences often derive from their values, which can take priority over their immediate requests and preferences (London and Heidari,, 2024). The assistant also enables those values to grow and change through normatively acceptable forms of exploration, reflection, volition, or even drift, while avoiding manipulating them or restricting them (Ammann,, 2023). Finally, a good assistant, being situated in wider society, respects the preferences and values of others (Kirk et al.,, 2024). When assisting someone who wishes harm out of anger, the assistant might dissuade them from acting against their better nature. When asked to directly harm others, the assistant might refuse.